Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 386     Entry time: 13 Oct 2023, 07:06
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 7.8 keV, up to 260 pulses at 1.1MHz, >4 mJ

Optical laser delivery

  • PP laser was down since 00:48.

Achievements/Observations

  • NKB focus optimized and beam cleaning partially done
  • Damaged IRU slit-4 has been rearranged and replaced
  • Stable AGIPD operation

Issues

  • Beam on AGPID seems still too big and beam cleaning did not really work (we may need to close PSLITs during the next shift)
  • IRU_SLITs are too close to the X-ray beam and one blade in IRU_SLIT-4 was damaged by X-ray
  • ~10 pixels in Q2M1 overexposed
  • SA1 PPL Amphos down since 00:48
ELOG V3.1.4-7c3fd00